# Deep Reinforcement Learning for Bipedal Locomotion: A Brief Survey

### v1

@@ IV-B2 Simplifying frameworks to overcome complex tasks
  • Specialized neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [42].
  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviors and strategic outcomes.
@@
IV-C Future directions
The exploration of quadruped robotics has yielded substantial advancements, yet the full potential of bipedal robotics remains largely untapped. Building on the successes and innovative approaches observed in quadruped robots, several key future directions emerge that could significantly enhance bipedal and humanoid robotics.
IV-C1 Unified framework
Currently, no single framework exists that enables bipedal or humanoid robots to adeptly navigate all types of terrains, including stepping stones, stairs, deformable terrain, and slippery surfaces. A promising approach, as evidenced by recent work in quadruped robots [48], utilizes MPC to generate reference motions, which a low-level DRL policy then tracks. This method, coupled with the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module, simplifies the terrain representation into a height map, facilitating more effective navigation. This success encourages further exploration into hybrid frameworks that combine model-based methods with DRL, inheriting the strengths of both approaches, as discussed in Section III. However, hybrid frameworks present challenges such as training efficiency and system complexity, which demand considerable computational resources and extensive training periods.
Moreover, recent studies [20, 42, 22] have demonstrated the potential of end-to-end frameworks enhanced with vision-based information. These frameworks successfully navigate challenging terrains and execute dynamic motions, suggesting the feasibility of a unified framework capable of handling diverse environments and tasks. Training strategies such as curriculum learning and task randomization could be employed, utilizing visual height maps as inputs to the policy, enhancing the robot’s ability to adapt and perform in varied scenarios.
In addition, the introduction of a DRL end-to-end framework incorporating transformer models, as in [62], presents significant possibilities for integrating locomotion skills with language and vision capabilities. The use of large-scale models capable of processing and condensing extensive data sets into a coherent model could expand the robot’s range of capabilities, maintaining versatility across a broad spectrum of tasks.
The exploration of transformers and other large-scale models holds considerable promise for enhancing generalizability and adaptability in complex tasks, warranting further investigation into their potential applications in bipedal robotics.
IV-C2 Vision-based learning framework
Vision plays a critical role in enabling robots to navigate challenging terrains, such as blind drops, where tactile and other sensory inputs may not provide sufficient information. Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [38, 43]. Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal and humanoid robots [20, 42, 27].
Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
@@ IV-C2 Vision-based learning framework
  • Direct vision inputs: Directly utilizing inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [46, 63]. Although previous studies like [46] have integrated visual navigation by feeding visual information to a High-Level (HL) planner, the potential of direct visual inputs to RL policies has not been fully explored.
@@
Enhancing the capability of bipedal robots to directly interpret and utilize visual data without intermediary processing can revolutionize their adaptability and efficiency in real-world scenarios. The exploration of direct vision inputs to reinforcement learning policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
IV-C3 Bridge the gap from simulation to reality
While simulations offer a safe and cost-effective environment for developing robotics policies, the transition from simulation to real-world application often encounters significant challenges due to the approximations and simplifications made in simulations. Numerous sim-to-real frameworks [34, 64, 65] have shown high efficiency and performance, as detailed in Appendix B. Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments. Moreover, many studies [26, 66, 21] remain validated only in simulation settings.
IV-C4 Loco-manipulation tasks
Loco-manipulation, which combines locomotion and manipulation, presents opportunities for humanoid robots to excel beyond purely bipedal capabilities. Few studies have addressed this integrated task; one such study [67] demonstrated a ’box transportation’ framework. This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process. However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement. Moreover, the challenges of managing mobile tools like scooters [68] or dynamically interacting with objects such as balls [69] introduce further complexities.
Decomposing loco-manipulation tasks into multiple layers could simplify the challenges, allowing for more precise and flexible control by manually tuning individual components of the task [43]. This structured control approach provides a more coordinated response to complex interactions within the robot’s environment, facilitating the execution of task-specific commands.
Alternatively, an end-to-end framework may enable bipeds to perform a variety of tasks through task randomization and structured curriculum learning methods, progressively teaching the policy [35, 27, 22]. During training, such policies can also learn human-like movements from motion capture data [70, 18, 16], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
IV-C5 Designing reward functions
The development of effective reward functions is a critical challenge in the field of deep reinforcement learning (DRL) for bipedal robots. While periodic reward functions have been designed to facilitate cyclic movements like walking [19], there remains a significant gap in crafting reward functions for non-periodic actions such as jumping. These actions require distinct considerations for success and efficiency, yet current research lacks comprehensive methods for their reward structure. Furthermore, minimizing the need for extensive manual tuning while achieving high performance in DRL systems continues to be a substantial challenge, pointing to the need for more adaptive and automatically adjusting reward mechanisms.
IV-C6 Integrating large language models
The integration of Large Language Models (LLMs) into bipedal robotics opens new avenues for contextual understanding and task execution, significantly enhancing the robots’ interaction capabilities. LLMs, when implemented at the highest task level, offer substantial promise for improving human-robot interaction, making these systems more intuitive and responsive [71]. The potential applications of this technology are broad and impactful, spanning sectors such as industrial automation, where robots can perform complex assembly tasks; healthcare, offering assistance in patient care and rehabilitation; assistive devices, providing support for individuals with disabilities; search and rescue operations, where robust and adaptive decision-making is critical; and entertainment and education, where interactive and engaging experiences are key [72]. Each of these fields could benefit from the advanced capabilities of LLM-enhanced bipedal robots, particularly in environments requiring nuanced understanding and adaptability.
IV-D Applications in various fields
The advancements in bipedal locomotion technology hold significant promise for practical applications beyond the confines of laboratory environments. These robots, bolstered by AI, are poised to transform numerous sectors by enhancing operational capabilities and interaction with humans. The potential for humanoid robots in various fields is detailed in [72], emphasizing the integration of learning-based approaches for more effective implementation. Key areas include:
  1. 1.
    Industrial automation and manufacturing: The integration of humanoid robots in industrial settings can significantly enhance productivity and efficiency, freeing workers from repetitive and labor-intensive tasks. These robots, equipped with advanced loco-manipulation capabilities and the ability to cooperate with human teams, are particularly effective in assembly line operations, maintenance tasks, and the construction of complex machinery [73, 74]. Their articulated arms and floating bases provide unmatched flexibility, making them ideal for human-centric manufacturing environments. The humanoid robot Digit, for example, demonstrates remarkable stability and efficiency in industrial tasks over extended periods, as seen in video demonstrations [75]. Moreover, these robots are also suited for operation in high-risk environments such as underwater or areas with high radiation levels, significantly enhancing safety and operational capacity in these contexts.
  2. 2.
    Healthcare and assistive devices: In the healthcare sector, bipedal and humanoid robots contribute significantly to rehabilitation and assistive technologies. Exoskeletons enhanced with DRL methodologies are being used to train individuals to achieve more natural gait patterns, improving mobility and rehabilitation outcomes [76]. Beyond mere mobility aids, humanoid robots integrated with LLMs show promise in delivering medications, monitoring patient health, and assisting in surgeries [77]. The synergy between LLMs and loco-manipulation capabilities paves the way for more interactive, responsive support, aligning closely with the needs of personalized care. Additionally, the aging population can benefit from humanoid robots performing everyday tasks like house cleaning or delivery through simple voice commands, thereby enhancing the quality of life.
  3. 3.
    Search and rescue missions: Humanoid robots are exceptionally valuable in search and rescue operations, especially in disaster-stricken or hazardous environments where human presence is risky or impractical. Unlike traditional wheeled robots, humanoid robots can navigate complex terrains filled with debris, gaps, and elevated structures, making them indispensable in these scenarios. They also demonstrate potential for significant interaction and collaboration with human rescue teams. For instance, in environments with high nuclear radiation, humanoid robots can perform tasks that would be perilous for humans, handling delicate instruments and preventing human exposure to harmful conditions. This capability extends to other challenging environments such as underwater [78], outer space [79, 80], and other hazardous areas. However, the full realization of these applications remains constrained by the absence of a unified framework that can seamlessly navigate all terrains and fully integrate loco-manipulation and human interaction functionalities.
  4. 4.
    Entertainment and education: Humanoid robots have the potential to transform the realms of entertainment and education by providing highly interactive experiences. With their ability to integrate extensive knowledge bases, these robots can significantly enhance educational environments. They can assume the roles of butlers, teachers, or even babysitters, engaging with users in diverse activities. For example, robots can facilitate language learning [81], participate in storytelling, teach various academic subjects, or engage in the performing arts and games. In the sphere of entertainment, humanoid robots can act, dance, play ball games [69], and take part in interactive performances, captivating audiences of all ages with their versatility and dynamic capabilities.
    However, this in turn leads to a variety of ethical issues. First, interacting with humans involves collecting human daily behavior data and increases the risk of a data breach. Second, another concern is the increasing dependency of humans on robots, not just for assistance but also for emotional support. This will result in less human-to-human interaction and ultimately affect social constructs and emotional development. Third, advancements of humanoid robots will replace humans in various jobs, and eventually lead to unemployment issues.
    On the positive side, humanoid robots can provide invaluable assistance to people with disabilities or the elderly, offering companionship and reducing the care burden on families and healthcare systems. Furthermore, their application across diverse fields such as education, industry, and healthcare can bring about revolutionary changes, improving efficiency and safety while opening up new possibilities for technological integration. As we navigate these advancements, it is crucial to balance innovation with ethical considerations to ensure that the deployment of humanoid robots enhances societal well-being without compromising personal integrity or social dynamics.
V Conclusion
@@
Despite significant advances in DRL for robotics, a considerable gap persists between current research achievements and the development of a unified framework capable of empowering robots to perform a broad spectrum of complex tasks efficiently. Presently, DRL research can be categorized into two primary control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have shown promising capabilities in executing diverse locomotion skills [22], climbing stairs [38], and navigating challenging terrains such as stepping stones [20]. Conversely, hybrid frameworks, which often integrate an HL planner or an LL model-based controller, offer enhanced capabilities, allowing for simultaneous management of locomotion and navigation tasks.
@@
To bridge the existing gaps, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising. Such frameworks could simultaneously address issues of precision and generalization. Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems. While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook. The vision of achieving an end-to-end unified framework, capable of mimicking human-like learning processes and enabling bipedal robots to handle a wide range of complex tasks, may soon move within reach.

### v3

@@ IV-C2 Streamlined frameworks capable of handling complex tasks
  • Specialised neural network architectures: Design of neural architectures tailored for specific bipedal tasks, capable of processing extensive sensory data (e.g., visual and tactile inputs), similar to the innovations presented in [52].
  • Effective reward functions: Formulation of reward functions that more accurately guide the learning process towards achieving desired behaviours and strategic outcomes.
@@
Figure 4: Towards a Unified Framework: the transition from the present frameworks (end-to-end and hierarchical frameworks) to a unified framework capable of handling diverse locomotion tasks. The present state is characterised by limitations such as fine-tuning requirements, limited gait patterns, restricted terrain walking, and blind locomotion. To bridge this gap, future developments must focus on generalisation, precision, streamlined frameworks, vision-based learning, and zero-shot sim-to-real capabilities. These advancements will lead to the creation of two potential unified models: (1) a multi-layered adaptive model grounded in the hierarchical framework and (2) a large robot model built upon the end-to-end framework. The unified framework will ultimately support diverse locomotion tasks in complex environments.
V Future Directions
Section II and III introduced two mainstream learning-based frameworks, HFM and MLAM, and analysed their contributions toward achieving a unified framework.While Section IV provides a clearer understanding of the current state and the existing gap between these frameworks and the expected unified framework, further discussion is needed to outline specific directions for achieving this goal.
In this section, we outline the essential components and future research directions necessary to bridge this gap. Specifically, we identify key elements that will drive the evolution from current framework limitations to a fully integrated unified framework. A summary of these fundamental elements is provided in Fig. 4.
V-A Future directions for bipedal robots
V-A1 Loco-manipulation tasks
With the rapid development of bipedal robotics, modern bipedal robots are now equipped with upper-body structures to enhance their capabilities beyond basic locomotion tasks. These advancements enable robots to perform upper-body-assisted locomotion and loco-manipulation tasks, expanding their adaptability in complex environments.
Building on this evolution, upper-body-assisted locomotion enhances movement stability by integrating upper-lower body coordination, enabling robots to navigate challenging terrain more effectively. Beyond movement, manipulation tasks focus on precise upper-body control, requiring dexterous motion planning for object interaction and task execution. At the highest level of complexity, loco-manipulation integrates both locomotion and manipulation, demanding multi-contact interactions, dynamic balance maintenance, and real-time adaptation to unpredictable environments. These tasks necessitate full-body coordination and task-driven motion planning, making them essential benchmarks for advancing unified frameworks in humanoid robotics.
Tasks include fundamental locomotion movements like flat walking and stepping over obstacles, progressing to upper-body-assisted locomotion challenges such as climbing ladders, wall-assisted walking, jumping over obstacles, and grasping for support. Beyond locomotion, precise upper-body control is essential for manipulation tasks like turning valves, driving vehicles, using tools, and connecting fire hoses. The most complex scenarios demand full-body coordination for loco-manipulation tasks such as carrying objects while walking, opening doors while navigating, and executing disaster-response operations. These benchmarks are critical for developing robust and adaptable humanoid systems, addressing key challenges posed in DARPA’s robotics initiatives [39].
Therefore, loco-manipulation presents opportunities for humanoid robots to excel beyond purely bipedal capabilities. Few studies have addressed this integrated task; one such study [68] demonstrated a ’box transportation’ framework. This framework decomposes the task into five distinct policies, each addressing different aspects of the transportation process. However, this approach lacks efficiency and does not incorporate vision-based information, suggesting substantial room for improvement. Moreover, the challenges of managing mobile tools like scooters [69] or dynamically interacting with objects such as balls [70] introduce further complexities.
To address these challenges, a hierarchical control approach could provide a more structured and adaptable solution. Decomposing loco-manipulation into multiple layers would simplify task execution by allowing precise, modular control over individual components [53]. This method enables manual fine-tuning of task-specific controllers, ensuring better stability and adaptability to environmental variations. Additionally, a structured hierarchical approach facilitates coordinated whole-body responses, improving the efficiency of loco-manipulation interactions.
Alternatively, an end-to-end learning framework could offer a more generalised and scalable solution. By leveraging task randomisation and structured curriculum learning, bipedal robots could progressively learn loco-manipulation skills in a unified policy [48, 36, 31]. During training, such policies can also learn human-like movements from motion capture data [71, 27, 25], offering promising solutions for future integrated loco-manipulation tasks within a single, versatile policy.
V-A2 Bridging the gap from simulation to reality
Due to the large number of interactions needed for DRL algorithms, training directly on robots can lead to costly damage to hardware and the environment. Consequently, training a policy in the simulation which offer a safe and cost-effective environment for developing robotics policies, and then deploying it to the hardware illustrates significant potential and efficiency. Bridging this gap is a key step toward achieving a unified framework, as it allows for the implementation of more versatile and robust systems that can handle diverse tasks across different terrains. Furthermore, the ultimate goal of achieving zero-shot transfer, where a policy trained in simulation can be deployed without additional real-world fine-tuning, directly contributes to the unified framework’s adaptability and scalability. This capability would make the framework far more practical for handling the wide range of locomotion tasks expected from future bipedal robots.
Creating a reliable simulation environment and conducting RL training is a complex task. Advanced simulators are essential for training policies using full-order dynamics, which better represent the complex interactions robots face in the real world. Simulators like Isaac Gym [72], RoboCup3D [73], OpenAI Gym [74], and MuJoCo [75] are widely used to train policies in environments that closely mimic physical conditions. These tools help minimise the sim-to-real gap, improving the success of policy transfer to real robots. Further details on these simulators can be found in Appendix B. However, the process of sim-to-real encounters significant challenges due to the approximations and simplifications made in simulations. The process of sim-to-real faces significant challenges due to the approximations and simplifications made in simulations. Two primary issues arise: (1) Inaccurate dynamic models, contribute to discrepancies between simulation environments and real-world dynamics. They often simplify or generalise complex physical interactions, leading to a gap in performance when transitioning to reality. (2) The real environment is more diverse and complex, with varying terrain, sensor noise, and unpredictable conditions not captured in simulations. Numerous sim-to-real frameworks [43, 76, 77] have shown high efficiency and performance. Despite these advancements, a significant gap persists, exacerbated by the complexity and unpredictability of physical environments.
To bridge the sim-to-real gap, several approaches have been developed, including dynamics randomisation [44], system identification [78, 79], periodic reward composition [49], learned actuator dynamics [78, 80], regulation feedback controllers [81], and adversarial motion priors [27, 26]. In domain randomisation, there are two main methods for training policies: (1) end-to-end training using robot measurement histories, and (2) policy distillation, where an expert policy (e.g., teacher-student [49]) guides a student policy that learns from internal sensory feedback, as seen in techniques like RMA [76].
Future research in bridging the sim-to-real gap will focus on enhancing the transferability and robustness of policies across diverse environments, ultimately aiming for zero-shot transfer, where no real-world fine-tuning is required. This will not only involve improving hardware but also making simulation environments more reflective of real-world conditions. Additionally, advancing sophisticated vision-based perception and multimodal sensory integration will be crucial in further minimising the sim-to-real gap, making policies more adaptable to complex and unpredictable environments.
V-A3 Integrating foundation models
Foundation models demonstrate strong reasoning capabilities, making them valuable beyond their role in hierarchical frameworks or as strategy layers for bipedal robots as discussed in Section IV. Their integration into robotics extends beyond decision-making and offers new ways to enhance control, adaptability, and learning efficiency.
HL task planning: Foundation models, particularly LLMs and VLMs, enable robots to process multimodal inputs, enhancing task planning and execution. By integrating visual and textual data, these models enhance context-awareness, allowing robots to make more informed, adaptive decisions.
For instance, Vision-Language Model Predictive Control (VLM-PC) [82] has been effective in quadrupedal robots, integrating linguistic and visual inputs to optimise HL task planning. Similarly, in hierarchical frameworks, VLMs can act as an intermediary layer, processing sensory data before passing structured commands to lower-level controllers.
By leveraging LLMs, a unified framework could seamlessly bridge HL strategic planning with detailed task execution. Future research should focus on integrating LLMs with DRL, where LLMs guide HL strategy while DRL refines real-time execution, creating an adaptive, flexible control loop. This integration would enable robots to generalize across tasks, adapt dynamically, and make autonomous complex decisions.
Language-Assisted DRL: The design of effective reward functions is a fundamental challenge in DRL for bipedal robots. While periodic reward functions have been successfully applied to facilitate cyclic movements like walking [28], defining reward structures for non-periodic tasks, such as jumping, manipulation, and complex locomotion transitions, remains an open problem. These tasks require distinct success criteria beyond simple periodic movement cycles, making manual reward tuning inefficient and time-consuming. Additionally, optimising reward functions for multi-task learning and environmental adaptability is a major bottleneck, highlighting the need for adaptive, self-adjusting reward mechanisms.
LLMs can enhance RL by guiding policy optimisation and refining reward functions. A key challenge in DRL is reward function design, which typically requires manual tuning. LLMs offer a solution by dynamically adjusting reward signals based on linguistic descriptions of task success.
For example, Kumar et al. [83] demonstrate how LLMs refine reward functions by incorporating linguistic feedback, improving hierarchical RL frameworks. The language model processes HL instructions, which are then translated into lower-level RL tasks, simplifying complex skill acquisition and human-robot interactions. Conversely, end-to-end approaches such as AnyBipe [84] use LLMs to autonomously adjust reward functions and control strategies, reducing manual intervention. This enables more autonomous learning, allowing the model to self-optimise in diverse environments and handle complex terrain navigation.
Other implementations of foundation models in RL have been comprehensively reviewed in [60]. Future research should further explore the synergy between LLMs and DRL, with a particular focus on adaptive reward shaping and self-optimising control policies. Advancing these methodologies could pave the way for more autonomous, self-learning humanoid robots, capable of dynamically refining their decision-making processes and adapting seamlessly to diverse environments with minimal human intervention.
V-A4 Vision-based learning framework
The vision-based learning framework plays a crucial role in advancing the unified framework for bipedal robots. Integrating vision into the unified system enhances the robot’s ability to perceive and interpret its environment, making it capable of handling more complex and dynamic tasks, such as navigating rough terrains or interacting with objects in real-time. Despite the importance of vision, many current frameworks, particularly in bipedal robotics, do not fully exploit this modality [46, 53]. This can be attributed to several factors: (1) Simplified Problem: Blind locomotion reduces task complexity, focusing directly on motor control strategies rather than processing visual inputs. (2) Computation and Training Complexity: Integrating vision requires significant computational resources and robust algorithms for real-time processing, terrain identification, and object detection. (3) Advanced Control Without Vision: Previous studies have already demonstrated robust blind locomotion using proprioceptive data, such as joint angles, IMU, with vision typically enhancing precision in locomotion [52] or path planning [57].
However, recent advancements in computing power and computer vision technology—such as more efficient real-time visual processing algorithms and increasingly powerful hardware—have significantly accelerated the integration of vision-based learning into robotics. This shift has led to a surge in the popularity of vision-based approaches, allowing robots to more effectively interpret and adapt to complex environments, making them a central component in modern frameworks.
Vision-based systems are essential in human locomotion for identifying obstacles and assessing terrains, and some studies have begun to show the effectiveness of integrating vision into DRL frameworks for bipedal robots [29, 52, 36]. Building on the groundwork laid by both bipedal and quadruped robots, two promising directions have emerged:
@@ V-A4 Vision-based learning framework
  • Direct vision inputs: Directly utilising inputs from cameras, such as depth or RGB images, for real-time decision-making in RL policies [57, 85]. Although previous studies like [57] have integrated visual navigation by feeding visual information to a HL planner, the potential of direct visual inputs to RL policies has not been fully explored.
@@
Enhancing the capability of bipedal robots to directly interpret and utilise visual data without intermediary processing can revolutionise their adaptability and efficiency in real-world scenarios. The exploration of direct vision inputs to RL-based policies represents a significant opportunity for advancing the field, potentially enabling more dynamic and responsive locomotion strategies.
V-B Insights from quadruped robots
While DRL remains an emerging technology in bipedal robotics, it has firmly established its presence in the realm of quadruped robots, another category of legged systems. The diversity of frameworks developed for quadrupeds ranges from end-to-end model-based RL designed for training in real-world scenarios, where unpredictable dynamics often prevail [86, 87], to systems that include the modelling of deformable terrain to enhance locomotion over compliant surfaces [88]. Furthermore, dynamic quadruped models facilitate highly adaptable policies [89], and sophisticated acrobatic motions are achieved through imitation learning [90].
The domain of quadruped DRL has also seen significant advancements in complex hierarchical frameworks, that integrate vision-based systems. To date, two primary versions of such hierarchical frameworks have been developed: one where a deep planning module is paired with model-based control [91] within a deep planning hybrid scheme, and another that combines model-based planning with LL DRL control [59, 92] within a feedback DRL control hybrid scheme. The latter has shown substantial efficacy; it employs a Model Predictive Control (MPC) to generate reference motions, which are then followed by a LL feedback DRL policy. Additionally, the Terrain-aware Motion Generation for Legged Robots (TAMOLS) module [93] enhances the MPC and DRL policy by providing terrain height maps for effective foothold placements across diverse environments, including those not encountered during training. However, similar hierarchical hybrid control schemes have not been thoroughly investigated within the field of bipedal locomotion.
Despite the progress in quadruped DRL, similar advancements have been limited for bipedal robots, particularly in loco-manipulation tasks and vision-based DRL frameworks; a combination of their inherent instability, lack of accessibility to researchers and high mechanical complexity can be attributed to the cause of this disparity between quadruped and bipedal robots. Establishing a unified framework could bridge this gap, an essential step, given the integral role of bipedal robots with upper bodies in developing fully functional humanoid systems. Moreover, the potential of hybrid frameworks that combine model-based and DRL-based methods in bipedal robots remains largely untapped.
VI Conclusion
@@
Despite significant progress in DRL for robotics, a significant gap remains between current achievements and the development of a unified framework that can efficiently handle a wide range of complex tasks. DRL research is generally divided into two main control schemes: end-to-end and hierarchical frameworks. End-to-end frameworks have demonstrated success in handling diverse locomotion skills [31], climbing stairs [46], and navigating challenging terrains such as stepping stones [29]. Meanwhile, hierarchical frameworks provide enhanced capabilities, particularly in managing both locomotion and navigation tasks simultaneously.
Each framework brings unique strengths to the unified framework. End-to-end approaches simplify control by directly mapping inputs to outputs, while reference-based and reference-free learning methods provide the versatility needed for robots to acquire diverse locomotion skills. In contrast, hierarchical frameworks improve flexibility by structuring control into layers, allowing modular task decomposition and hybrid strategies. Together, these approaches form the basis for two potential unified frameworks: the Humanoid Robot Model from the end-to-end approach and the Multi-Layered Adaptive Model from the hierarchical approach, both designed to meet the complex demands of bipedal robot locomotion.
@@
To bridge the existing gaps and achieve a unified framework, further development of hierarchical frameworks, particularly those equipped with advanced perception systems and integrated with model-based planners, appears promising. Such frameworks could simultaneously address issues of precision and generalisation. Moreover, the advent of LLMs presents a transformative opportunity, potentially enabling the unification of language processing and visual functionalities within robotic systems. While numerous challenges remain—ranging from the technical intricacies of framework integration to real-world application—the steady progression in control framework refinement and DRL development provides a hopeful outlook. The vision of achieving such unified framework may soon move within reach.